OpenStudy (anonymous):

Prove that if m and n are perfect squares then m + n +2 √ mn is also a perfect square. Using direct proof.

OpenStudy (blurbendy):

so what do you get to assume in a direct proof?

OpenStudy (blurbendy):

hint: p -> q

OpenStudy (anonymous):

I assume that n is a perfect square

OpenStudy (blurbendy):

m and n are perfect squares

OpenStudy (blurbendy):

so if that's what we get to assume, it will be helpful if we can find a way to rewrite m and n

OpenStudy (blurbendy):

for example, an even number can be written as 2k, and an odd number can be written as 2k + 1 by definition. any ideas on how we can represent a perfect square?

OpenStudy (blurbendy):

Here's a hint: A perfect square is a number that can be expressed as the product of two equal integers.

OpenStudy (anonymous):

I'm actually excited to see this one. I have an idea of what I would use, but I don't want to hijack the OP and rob him of his eureka moment.

OpenStudy (blurbendy):

hehe

OpenStudy (anonymous):

(n+2)^2

OpenStudy (blurbendy):

that is in the form of a perfect square, but i dont think it would be helpful in our proof. if a perfect square can be expressed as the product of two equal integers, then if we say x is some integer, then m = (x*x) = x^2 fits our definition pretty nicely

OpenStudy (blurbendy):

the same can be done for n now, so how would you write that

OpenStudy (anonymous):

n = (x*x) = x^2

OpenStudy (blurbendy):

hm, i dont think we get to assume that m and n are the same perfect square, so instead of x, let's choose y

OpenStudy (anonymous):

okay. n = (y*y) = y^2

OpenStudy (blurbendy):

okay, so now that we have definitions for m and n, we can apply them to what we are trying to prove: namely m + n + 2\[\sqrt{mn}\]

OpenStudy (blurbendy):

sqrt[mn] should be next to the 2, dont know why it moved down

OpenStudy (blurbendy):

go ahead and substitute for m and n

OpenStudy (blurbendy):

actually, now that i think about it, i think m and n are going to have to be equal.

OpenStudy (blurbendy):

but go ahead and continue

OpenStudy (anonymous):

Are we sure about not being able to assume they can use the same bound variable? I'm really new to this also, but it looks to me like it works out if we just use x^2 for m and x^2 for n.

OpenStudy (blurbendy):

yeah, they will have to be, because in order for our proof to work in the end, m + n + 2sqrt[mn] will need to be a perfect square, and by our definition, a perfect square can be expressed as the product of two equal integers

OpenStudy (blurbendy):

do you know what the next step is @mtalampas21 ?

OpenStudy (anonymous):

I'm working my way through Vellman's How to Prove It and I thought I remembered something about even with different free variables, it isn't a guarantee that different object will use them (without a y doens't equal x provision).

OpenStudy (anonymous):

I'm not really sure

OpenStudy (blurbendy):

interesting @Ajk. I read like 1/3 of that of that book, then i got too busy. interesting read from what i gathered

OpenStudy (blurbendy):

okay, we are going to let m and n both = x^2, and ill show you why afterwards. since we have those definitions, we can use substitution, for what we are trying to prove. so wherever you see m and n, put x^2 in there place

OpenStudy (blurbendy):

their*

OpenStudy (anonymous):

I'm liking the next step being just filling in m and n with what @blurbendy suggested (x^2) and doing the math. I ended up with a perfect square once it was all added together in the end.

OpenStudy (anonymous):

x^2 + x^2 + 2 sqrt(x^2)(x^2)

OpenStudy (anonymous):

I don't know if I'm right, but that is what I did. Once you perform the root operation and add everything up . . .

OpenStudy (blurbendy):

okay so after some simplifying you should get 2x^2 + 2sqrt[x^4]

OpenStudy (blurbendy):

what is the square root of x^4?

OpenStudy (blurbendy):

dont give up, you're close!

OpenStudy (anonymous):

x^2

OpenStudy (blurbendy):

good, so that gives us 2x^2 + 2x^2. looking good, we need it in the form of a perfect square, and a perfect square can be expressed as the product of two equal integers. 2x^2 + 2x^2 gives us 4x^2 Now, we can we rewrite 4x^2 so that it resembles the definition of a perfect square?

OpenStudy (blurbendy):

how can we **

OpenStudy (anonymous):

2x * 2x

OpenStudy (blurbendy):

not quite

OpenStudy (blurbendy):

think of another way to write 4

OpenStudy (blurbendy):

actually, that might work

OpenStudy (anonymous):

4(x*x)

OpenStudy (blurbendy):

that will work, i just did it differently

OpenStudy (anonymous):

yeah, I think any. (2x)^2 2^2*x^2 2x*2x . This was a good question. Thanks guys.

OpenStudy (blurbendy):

yup, good job mtalampas. hope it makes sense. do we we see what we concluded though? 2x * 2x = the product of two equal integers = a perfect square which is what we were trying to show.

OpenStudy (anonymous):

yup, it makes sense. Thanks a lot.

OpenStudy (blurbendy):

you bet.

OpenStudy (anonymous):

may you give me another example of a direct proof? If it's okay

OpenStudy (blurbendy):

let me think.... okay Show that the product of two rational numbers is rational.

OpenStudy (blurbendy):

We get to assume that we have two rational numbers, and we to show their product is rational

OpenStudy (blurbendy):

similar to what we did for a perfect square, we need some sort of definition to rewrite those numbers. Let's call them m and n again

OpenStudy (blurbendy):

any ideas on how to write a rational number?

OpenStudy (anonymous):

m/n, where n is not equal to zero ?

OpenStudy (blurbendy):

perfect, and we also need to say a/b are in the set of integers (since we dont want fractions). also instead of m/n, let's let m = a/b, where b is not zero and a and b are in the set of integers and n = x /y where y is not zero and x and y are in the set of integers

OpenStudy (blurbendy):

so now that we have our definitions, we can substitute them in for m*n, and then we'll show that their product is rational

OpenStudy (blurbendy):

so, what will our product look like

OpenStudy (anonymous):

a/b * x/y

OpenStudy (blurbendy):

yup, and that gives us (ax) / (by), and notice we have it in the form p / q where q isn't 0 and p and q are in the set of integers. this the definition of rational, so we have concluded that the product of two rational numbers is always rational

OpenStudy (anonymous):

I get it, thanks.

OpenStudy (blurbendy):

Direct proofs are pretty straight forward. Label your assumptions. State a way to rewrite your assumptions according to some definition. apply your rewritten assumptions to what you are trying to prove, and the math should take care of itself.

OpenStudy (anonymous):

How about Indirect proofs, by contrapositive and contradiction? Any examples.

OpenStudy (blurbendy):

for contrapositive and contradiction you have to be careful of what you get to assume. for example if you wanted to prove using contraposition that: if the product m*n is even, then m and n cannot both me odd in contraposition, you get to assume the negation of q and you have to show the negation of p so, in the above, you would get to assume m and n are both EVEN, and would have to show m*n is ODD In contradiction, you get to assume p and the negation of q so if you had: There does not exist an integer x, such that x^2 + 1 is negative You get to assume some integer x, and that x^2 + 1 is POSITIVE In contradiction, you go until you break math So if we are assuming there does not exist an integer x such that x^2 + 1 is POSITIVE, let's plug in x = 1, and show that that is true (1)^2 + 1 = 2 which is positive We reached a contradiction! Therefore, we can restate that our conjecture is true (There does not exist an integer x, such that x^2 + 1 is negative)

OpenStudy (blurbendy):

an indirect proof is the same as contradiction

OpenStudy (blurbendy):

well sort of

OpenStudy (blurbendy):

you basically show that it can't be false

OpenStudy (blurbendy):

i.e. the math never breaks

OpenStudy (anonymous):

both contradiction and contraposition are indirect proofs if I'm not mistaken..

OpenStudy (blurbendy):

probably. you dont see many proofs using the indirect method solely. most rely on contraposition or contradiction instead.

OpenStudy (anonymous):

I see. How about the counter proof?

OpenStudy (blurbendy):

these are the easiest in my opinion. they are mostly used as disproofs so if you have a statement that said There does not exist an integer x such that x^2 -1 is negative we can do a proof by counter-example since x is an integer, we can let x = 0 (0)^2 - 1 = -1 which is negative. hence, the statement "There does not exist an integer x such that x^2 -1 is negative" is false.

OpenStudy (blurbendy):

disproof by counter-example *** didnt mean to say proof

OpenStudy (anonymous):

I understand...Thank you =)

OpenStudy (blurbendy):

anytime. proofs are all about practice. do a bunch of em' and soon enough they'll become second nature.

OpenStudy (anonymous):

Noted. I'm so glad that I found this site, you guyz are really helpful. May I know your age and profession?

OpenStudy (blurbendy):

I'm 21, and i'm going to school for computer science right now. i'm actually interning for OpenStudy this summer

OpenStudy (anonymous):

Okay. I'm just curious. I'll go practice making proofs.

OpenStudy (blurbendy):

are you in highscool, college?

OpenStudy (blurbendy):

highschool*

OpenStudy (anonymous):

2nd year college. I'm 17

OpenStudy (anonymous):

I'm not good in math that's why.

OpenStudy (blurbendy):

cool, you'll be done early then

OpenStudy (blurbendy):

is this class part of your major?

OpenStudy (anonymous):

yup, in Discrete math. I'm an Information technology student

OpenStudy (blurbendy):

oh, right on

OpenStudy (anonymous):

in what country are you by the way?

OpenStudy (blurbendy):

United States, you?

OpenStudy (anonymous):

Oh. Philippines

OpenStudy (blurbendy):

ah, I have a friend from there. it's getting late though, so i'm going to take off. hope to see you around here again!

OpenStudy (anonymous):

Thanks again. This is not my last in this site for sure. =)

OpenStudy (blurbendy):

awesome. cya!